# Generalizable Quantum Computing Pipeline for Real World Drug Discovery

### v1

@@ Abstract
Quantum computing, with its superior computational capabilities compared to classical approaches, holds the potential to revolutionize numerous scientific domains, including pharmaceuticals. However, the application of quantum computing for drug discovery has primarily been limited to proof-of-concept studies, which often fail to capture the intricacies of real-world drug development challenges. In this study, we diverge from conventional investigations by developing an advanced quantum computing pipeline tailored to address genuine drug design problems. Our approach underscores the pragmatic application of quantum computation and propels it towards practical industrial adoption. We specifically construct our versatile quantum computing pipeline to address two critical tasks in drug discovery: the precise determination of Gibbs free energy profiles for prodrug activation involving covalent bond cleavage, and the accurate simulation of covalent bond interactions. This work serves as a pioneering effort in benchmarking quantum computing against veritable scenarios encountered in drug design, especially the covalent bonding issue present in both of the case studies, thereby transitioning from theoretical models to tangible applications. Our results demonstrate the potential of a quantum computing pipeline for integration into real world drug design workflows.
I Introduction
@@
Quantum computing is emerging as a powerful change that promises to significantly enhance scientific computing and simulations. Quantum computers, operating with quantum bits (qubits), have the potential to execute complex calculations at speed and levels of precision that traditional supercomputers cannot achieve [1, 2, 3]. The realm of drug discovery, characterized by its need for meticulous molecular modeling and predictive analytics [4, 5, 6, 7], stands as an ideal candidate to benefit from this quantum leap. Recent endeavors have commenced the integration of quantum computing into drug design research, marking a progressive stride in the application of advanced computational technologies to drug discovery [8, 9, 10]. In drug design, existing classical computational chemistry methods are not able to compute exact solutions, and the required computational cost grows exponentially as the scale of the system grows. Quantum algorithms exemplified by the Variational Quantum Eigensolver (VQE) [11], hold the potential to advance classical methods like Hartree-Fock (HF) [12] towards more accurate solutions within the quantum computing paradigm. As the scale of quantum computers expands, quantum computing approaches are expected to significantly outperform existing solutions, such as Density Functional Theory (DFT) [13], in terms of both accuracy and efficiency in scenarios involving quantum chemical calculations.
@@
However, in the current landscape, the involvement of quantum computing in drug discovery is primarily restricted to conceptual validation, with minimal integration into real world drug design. Our versatile quantum computing pipeline (see Figure 1) for real-world drug discovery addresses this gap by investigating two pertinent case studies rooted in actual clinical and pre-clinical contexts. The key step for quantum computation of molecular properties is to prepare the molecular wave function on a quantum device. To this end, the VQE framework is suitable for near-term quantum computers. As shown in Fig. 1(b), the core of VQE is to employ parameterized quantum circuits to measure the energy of the target molecular system. Then, a classical optimizer is employed to minimize the energy expectation until convergence. Due to the variational principle, the state of the quantum circuit becomes a good approximation for the wave function of the target molecule, and the measured energy is the variational ground state energy. After that, additional measurements can be performed on the optimized quantum circuit for other interested physical properties.
@@
Quantum computing can enhance our understanding of such drug-target interactions through QM/MM (Quantum Mechanics/Molecular Mechanics) simulations, which are vital in the post-drug-design computational validation phase. To realize this, we implemented a quantum computing workflow for molecular forces during QM/MM simulation. This development not only facilitates a detailed examination of covalent inhibitors like Sotorasib, but also propels the field of computational drug development forward.
Through these two real-world drug design examples, we present a versatile and unprecedented quantum computing pipeline for drug design. Our workflow has advantages in its flexibility and has been carefully constructed to accommodate various applications in the area of drug discovery. The universality of our pipeline highlights its potential as a foundational tool, empowering researchers with a ready-to-use computational resource.
@@
II Results
II.1 Gibbs Free Energy Profiles in Prodrug Activation Strategy
II.1.1 Carbon-carbon bond cleavage in prodrug activation strategy
@@
II.1.2 Gibbs free energy profiles of covalent bond cleavage for prodrug activation
@@
Despite that quantum devices with more than 100 qubits are becoming available, simulating large chemical systems would require very deep circuits, which will inevitably lead to inaccurate outcomes due to intrinsic quantum noise. Additionally, the $N^{4}$ terms to measure to calculate molecular energy is another bottleneck for quantum computation due to the limited measurement shot budget. Thus, it is often desirable to reduce the effective problem size of chemical systems, so that they can be processed on available quantum devices. The quantum embedding methods and downfolding methods have gathered significant attention recently [37, 38]. In this work, we employed the active space approximation due to its popularity and versatility, which simplifies the QM region into a more manageable 2 electron/2 orbital system. The fermionic Hamiltonian is then converted into a qubit Hamiltonian using parity transformation. The wave function of the active space can then be represented by a 2-qubit superconducting quantum device. We utilized a hardware-efficient $R_{y}$ ansatz with a single layer as the parameterized quantum circuit for VQE, as depicted in Fig. 3. We applied standard readout error mitigation to enhance the accuracy of the measurement results. For more detailed technical information, please refer to the Methods section. We implemented the entire workflow in the TenCirChem package [39], allowing users to utilize these functions with just a few lines of code.
@@
HF
CASCI
VQE
DFT
$\Delta G^{{\ddagger}}$ with solvent
13.1
11.8
$7\pm 5$
8.3
$\Delta G$ with solvent
-41.1
-49.5
$-51\pm 11$
-10.8
$\Delta G^{{\ddagger}}$ w/o solvent
2.6
1.4
$-4\pm 7$
$\Delta G$ w/o solvent
-59.3
-68.0
$-65\pm 9$
@@
Computational Wall Time (s)
Molecule CASCI VQE
4 358 424
5 3 63
6 97 161
TS 360 424
In this study, we have limited the utilization of quantum computers to a few qubits employing the active space approximation, due to the limited size and gate noise of currently available quantum computers. Herein, we estimate what kind of quantum computers are required for a fully correlated computation of the systems studied in this work without incurring the active space approximation. Taking molecule 4 as an example, with 6-311G(d,p) basis set, the system corresponds to $N=630$ orbitals and $N_{\textrm{elec}}=196$ electrons. To reduce the qubit requirement, the paired unitary coupled-cluster ansatz can be employed, which requires only 1 qubits for each orbital due to the restriction of electron pairing [42, 43]. Other advantages of the ansatz include that evaluating the energy requires only constant measurement and linear circuit depth due to the efficient Givens-SWAP network. Additionally, since there are $N_{\textrm{elec}}=196$ electrons, the number of all possible double excitations is $\frac{N_{\textrm{elec}}}{2}\times(N-\frac{N_{\textrm{elec}}}{2})=52136$. Thus, a fully correlated computation at PUCC level with double excitations (PUCCD) involves a quantum circuit with approximately $10^{3}$ qubits and $10^{5}$ Givens-SWAP gates.
@@
II.2 Covalent Bond Simulation
@@
II.2.1 KRAS and covalent inhibition
@@
Figure 5: (a): Energy evolution during the MD simulation. The energy stabilized after an initial equilibrating phase. (b): Monitoring the QM region energy evolution. (c) The covalent bond is remarkably stable during the whole simulation process. The bond length fluctuates around 1.86 angstrom with a standard deviation less than 0.1 angstrom. The bond length is aligned with previous literature discoveries, and the small deviation is a good indication of the bonding stability.(d) A visualisation of the bond angles (CB-SG-C18 and SG-C18-C17) and torsion angle (CB-SG-C18-C17) variation. The left plot is the time evolution of the two bond angles, the middle plot is their histograms, and the right plot is the time evolution of the dihedral angle. A high level of stability can be observed in all three, indicating that the inhibitor has been able to firmly bond to the G12C mutation through the covalent bond, promising high efficacy in its inhibition of KRAS signaling.
@@
II.2.2 QM energy update using quantum computers
@@
Considering the computational load, to check the soundness of the computation, we run the first 1600 steps of the simulation on a quantum computer as a sanity check, and the results closely follow the baseline QM/MM simulation, as can be observed in Fig 6. We then moved some key steps of the QM/MM simulation to the quantum computer, to establish a QM/MM-QC hybrid simulation system. In Fig 7a, the simulation is started on the quantum computer and continued on a classic computer; in Fig 7a, the simulation is started on a classical computer and continued on a quantum machine. Compared with the previous QM/MM simulation, we can see that the hybrid simulations have been able to closely follow the baseline trajectory, which gives us confidence that such hybrid simulations are a feasible use of the limited quantum computer computation powers.
@@
Method Total Time (minutes) Avg Time per Step (seconds)
Classic 124.0 4.65
Noiseless 126.0 4.725
Quantum 3820.0 143.25
@@
III Discussions
In this study, we have established a pipeline that enables quantum computers to tackle real-world drug discovery problems. Specifically, we have addressed two crucial challenges of computer-aided drug design, computing reaction barriers and molecular dynamics simulation. Our pipeline combines quantum-classical hybrid computing platforms, leveraging the VQE framework on the quantum computing side to efficiently store and manipulate molecular wave functions. On the classical computing side, we employ the ddCOSMO solvation model to compute solvation energy and analytical CASCI force formula to compute molecular forces for QM/MM simulation, respectively. The interface between the quantum and classical computing sides relies on the one and two-body reduced density matrices.
@@ III Discussions
Based on the two cases, we provide evidence that our quantum computing pipeline has the capability to contribute to solving real-world drug design problems. However, it is important to note that the accuracy of VQE calculations and the resources consumed require further improvement. On the quantum hardware side, continuous efforts should be made to enhance gate fidelity and strive toward achieving error correction. In terms of quantum algorithms, advanced VQE additions such as neural networks or Clifford circuits [51, 52] can be explored to enhance the accuracy of the VQE circuit. While our study employed classical pre-optimization instead of parameter optimization on quantum computers due to associated overhead, the next step in the development of the VQE pipeline should involve better circuit parameter initialization and more efficient parameter optimization algorithms. This will enable the complete transfer of the pipeline onto quantum computers, further leveraging their computational power.
The quantum computing methodologies developed in this study have the potential to extend beyond the presented case studies of Sotorasib and $\beta$-lapachone. The integration of quantum computing into QM/MM simulations offers a versatile platform that can be adapted and scaled to address a wide range of molecular targets and complex biological interactions.
IV Methods
IV.1 Quantum Computing for Molecular Systems
@@
The active-space approximation is employed to reduce computational cost and enhance accuracy. The approximation adopts the Hartree-Fock state as the baseline state and chooses an “active space” that is treated with a high-accuracy computational method. In classical computation, the high-accuracy method is usually full configuration interaction (FCI) or density matrix renormalization group (DMRG). In our case, quantum computers are employed to solve the problem with the VQE algorithm. The active space is usually constructed in the molecular orbital space. Most commonly, orbitals that have the closest energy with the HOMO and LUMO orbitals will be included in the active space. Meanwhile, the inner shell orbitals are treated at the mean-field level. Thus, this approximation is sometimes also called the frozen core approximation. Denote the set of frozen occupied spin-orbitals by $\Lambda$. The frozen core provides an effective repulsion potential $V^{\rm{eff}}$ to the remaining electrons
@@
IV.2 Quantum Computation of Solvation Effect

### v3

# A Hybrid Quantum Computing Pipeline for Real World Drug Discovery

@@ Abstract
Quantum computing, with its superior computational capabilities compared to classical approaches, holds the potential to revolutionize numerous scientific domains, including pharmaceuticals. However, the application of quantum computing for drug discovery has primarily been limited to proof-of-concept studies, which often fail to capture the intricacies of real-world drug development challenges. In this study, we diverge from conventional investigations by developing a hybrid quantum computing pipeline tailored to address genuine drug design problems. Our approach underscores the application of quantum computation in drug discovery and propels it towards more scalable system. We specifically construct our versatile quantum computing pipeline to address two critical tasks in drug discovery: the precise determination of Gibbs free energy profiles for prodrug activation involving covalent bond cleavage, and the accurate simulation of covalent bond interactions. This work serves as a pioneering effort in benchmarking quantum computing against veritable scenarios encountered in drug design, especially the covalent bonding issue present in both of the case studies, thereby transitioning from theoretical models to tangible applications. Our results demonstrate the potential of a quantum computing pipeline for integration into real world drug design workflows.
1 Introduction
@@
Quantum computing is emerging as a powerful change that promises to significantly enhance scientific computing and simulations. Quantum computers, operating with quantum bits (qubits), have the potential to execute complex calculations at speed and levels of precision that traditional supercomputers cannot achieve [1, 2, 3]. The realm of drug discovery, characterized by its need for meticulous molecular modeling and predictive analytics [4, 5, 6, 7], stands as an ideal candidate to benefit from this quantum leap. Recent endeavors have commenced the integration of quantum computing into drug design research, marking a progressive stride in the application of advanced computational technologies to drug discovery [8, 9, 10, 11]. In drug design, existing classical computational chemistry methods are not able to compute exact solutions, and the required computational cost grows exponentially as the scale of the system grows. Quantum algorithms exemplified by the Variational Quantum Eigensolver (VQE) [12], hold the potential to advance classical methods like Hartree-Fock (HF) [13] towards more accurate solutions within the quantum computing paradigm. As the scale of quantum computers expands, quantum computing approaches are expected to significantly outperform existing solutions, such as Density Functional Theory (DFT) [14], in terms of both accuracy and efficiency in scenarios involving quantum chemical calculations. In addition to the quantum chemistry approach, a variety of drug-design problems can be cast into optimization problems [15, 11]. The quantum approximate optimization algorithm [16] or quantum annealing algorithms [17, 18] can then be employed to solve these optimization algorithms.
@@
However, in the current landscape, the involvement of quantum computing in drug discovery is primarily restricted to conceptual validation, with minimal integration into real world drug design[19, 20, 21, 22, 23, 24]. Our hybrid quantum computing pipeline (see Figure 1) is real-world drug discovery problem oriented. Our approach addresses this gap by investigating two pertinent case studies rooted in actual clinical and pre-clinical contexts. The key step for quantum computation of molecular properties is to prepare the molecular wave function on a quantum device. To this end, the VQE framework is suitable for near-term quantum computers. As shown in Fig. 1(b), the core of VQE is to employ parameterized quantum circuits to measure the energy of the target molecular system. Then, a classical optimizer is employed to minimize the energy expectation until convergence. Due to the variational principle, the state of the quantum circuit becomes a good approximation for the wave function of the target molecule, and the measured energy is the variational ground state energy. After that, additional measurements can be performed on the optimized quantum circuit for other interested physical properties.
@@
Quantum computing can enhance our understanding of such drug-target interactions through QM/MM (Quantum Mechanics/Molecular Mechanics) simulations, which are vital in the post-drug-design computational validation phase. To realize this, we implemented a hybrid quantum computing workflow for molecular forces during QM/MM simulation. This development not only facilitates a detailed examination of covalent inhibitors like Sotorasib, but also propels the field of computational drug development forward.
Through these two real-world drug design examples, we present a hybrid quantum computing pipeline for drug design. Our workflow has advantages in its flexibility and has been carefully constructed to accommodate various applications in the area of drug discovery. The universality of our pipeline highlights its potential as a foundational tool, empowering researchers with a ready-to-use computational resource.
@@
2 Results
2.1 Gibbs Free Energy Profiles in Prodrug Activation Strategy
2.1.1 Carbon-carbon bond cleavage in prodrug activation strategy
@@
2.1.2 Gibbs free energy profiles of covalent bond cleavage for prodrug activation
@@
Despite that quantum devices with more than 100 qubits are becoming available, simulating large chemical systems would require very deep circuits, which will inevitably lead to inaccurate outcomes due to intrinsic quantum noise. Additionally, the $N^{4}$ terms to measure to calculate molecular energy is another bottleneck for quantum computation due to the limited measurement shot budget. Thus, it is often desirable to reduce the effective problem size of chemical systems, so that they can be processed on available quantum devices. The quantum embedding methods and downfolding methods have gathered significant attention recently [48, 49]. In this work, we employed the active space approximation due to its popularity and versatility, which simplifies the QM region into a more manageable 2 electron/2 orbital system. The CASCI energy can be considered as the exact solution under the active space approximation and the results by quantum computers are expected to be consistent with the CASCI energy. The fermionic Hamiltonian is then converted into a qubit Hamiltonian using parity transformation. The wave function of the active space can then be represented by a 2-qubit superconducting quantum device. We utilized a hardware-efficient $R_{y}$ ansatz with a single layer as the parameterized quantum circuit for VQE, as depicted in Fig. 3. We applied standard readout error mitigation to enhance the accuracy of the measurement results. For more detailed technical information, please refer to the Methods section. We implemented the entire workflow in the TenCirChem package [50], allowing users to utilize these functions with just a few lines of code.
@@
In this study, we have limited the utilization of quantum computers to a few qubits employing the active space approximation, due to the limited size and gate noise of currently available quantum computers. Herein, we estimate what kind of quantum computers are required for a fully correlated computation of the systems studied in this work without incurring the active space approximation. Taking molecule 4 as an example, with 6-311G(d,p) basis set, the system corresponds to $N=630$ orbitals and $N_{\textrm{elec}}=196$ electrons. To reduce the qubit requirement, the paired unitary coupled-cluster ansatz can be employed, which requires only 1 qubits for each orbital due to the restriction of electron pairing [53, 54]. Other advantages of the ansatz include that evaluating the energy requires only constant measurement and linear circuit depth due to the efficient Givens-SWAP network. Additionally, since there are $N_{\textrm{elec}}=196$ electrons, the number of all possible double excitations is $\frac{N_{\textrm{elec}}}{2}\times(N-\frac{N_{\textrm{elec}}}{2})=52136$. Thus, a fully correlated computation at PUCC level with double excitations (PUCCD) involves a quantum circuit with approximately $10^{3}$ qubits and $10^{5}$ Givens-SWAP gates. The PUCCD ansatz has been successfully implemented on both superconducting and trapped-ion quantum computers [53, 54]. The number of qubits employed in these studies is around 10, sufficient to describe 1 to 2 heavy atoms if active space or embedding techniques are not used [55]. While digital quantum computers with over 100 qubits are accessible [3], their application in quantum chemistry has been limited, primarily due to the restricted fidelity of two-qubit gates. However, with improved two-qubit gate fidelity, these quantum computers can handle complex molecules comprising dozens of atoms, such as molecule 4.
@@
2.2 Covalent Bond Simulation
@@
2.2.1 KRAS and covalent inhibition
@@
2.2.2 QM energy update using quantum computers
@@
Considering the computational load, to check the soundness of the computation, we run the first 1600 steps of the simulation on a quantum computer as a sanity check, and the results closely follow the baseline QM/MM simulation, as can be observed in Fig 6. We then moved some key steps of the QM/MM simulation to the quantum computer, to establish a QM/MM-QC hybrid simulation system. In Fig 7a, the simulation is started on the quantum computer and continued on a classic computer; in Fig 7b, the simulation is started on a classical computer, continued on a quantum machine, and subsequently moved back to the classic computer. Compared with the previous QM/MM simulation, we can see that the hybrid simulations have been able to closely follow the baseline trajectory, which gives us confidence that such hybrid simulations are a feasible use of the limited quantum computer computation powers.
@@
3 Discussions
In this study, we have established a model pipeline that enables quantum computers to tackle real-world drug discovery problems. Specifically, we have addressed two crucial challenges of computer-aided drug design, computing reaction barriers and molecular dynamics simulation. Our pipeline combines quantum-classical hybrid computing platforms, leveraging the VQE framework on the quantum computing side to efficiently store and manipulate molecular wave functions. On the classical computing side, we employ the ddCOSMO solvation model to compute solvation energy and analytical CASCI force formula to compute molecular forces for QM/MM simulation, respectively. The interface between the quantum and classical computing sides relies on the one and two-body reduced density matrices.
@@ 3 Discussions
Based on the two cases, we provide evidence that our hybrid quantum computing pipeline has the potential to solve real-world drug design problems. However, it is important to note that the accuracy of VQE calculations and the resources consumed require further improvement. On the quantum hardware side, continuous efforts should be made to enhance gate fidelity and strive toward achieving error correction. In terms of quantum algorithms, advanced VQE additions such as neural networks or Clifford circuits [63, 64] can be explored to enhance the accuracy of the VQE circuit. While our study employed classical pre-optimization instead of parameter optimization on quantum computers due to associated overhead, the next step in the development of the VQE pipeline should involve better circuit parameter initialization and more efficient parameter optimization algorithms. This will enable the complete transfer of the pipeline onto quantum computers, further leveraging their computational power.
While there are plenty of works in leveraging quantum computing for drug discovery[19, 20, 21, 22, 23, 24], the focus of our pipeline is for tackling specific real-world drug design problems. We emphasize the use of a convenient, modular, and hybrid quantum pipeline for drug discovery, which will make it more accessible for drug design experts without a quantum computing background. Additionally, referencing established criteria in the drug design domain, our computational results indicate that they also fall within reasonable bounds. Furthermore, the quantum computing methodologies developed in this study have the potential to extend beyond the presented case studies of Sotorasib and $\beta$-lapachone. The integration of quantum computing into QM/MM simulations offers a versatile platform that can be adapted and scaled to address a wide range of molecular targets and complex biological interactions.
4 Methods
4.1 Quantum Computing for Molecular Systems
@@
The active-space approximation is employed to reduce computational cost and enhance accuracy. The approximation adopts the Hartree-Fock state as the baseline state and chooses an “active space” that is treated with a high-accuracy computational method. In classical computation, the high-accuracy method is usually full configuration interaction (FCI) or density matrix renormalization group (DMRG) [65]. In our case, quantum computers are employed to solve the problem with the VQE algorithm. The active space is usually constructed in the molecular orbital space. Most commonly, orbitals that have the closest energy with the HOMO and LUMO orbitals will be included in the active space. Meanwhile, the inner shell orbitals are treated at the mean-field level. Thus, this approximation is sometimes also called the frozen core approximation. Denote the set of frozen occupied spin-orbitals by $\Lambda$. The frozen core provides an effective repulsion potential $V^{\rm{eff}}$ to the remaining electrons
@@
4.2 Quantum Computation of Solvation Effect
@@
4.3 Quantum Computation of Molecular Forces
[…]
4.4 Quantum Computation Details
$L_{\rm{CNOT}}^{(l)}$ $=\prod_{j=N-1}^{1}\textrm{CNOT}[j,j+1],$ (12)
$L_{R_{y}}^{(l)}(\bm{\theta})$ $=\prod_{j=N}^{1}R_{y}[{j}](\theta_{lj}).$
4.5 Classical Computation Details
4.5.1 Methods for obtaining the optimal geometry configuration
4.5.2 System preparation for covalent bond simulation
Our simulation started with the intricate system preparation using Amber software suite [86], especially packages including pdb4amber, antechamber, parmchk2, tleap, etc, defining the fundamental molecular and environmental parameters. This initial setup was crucial in modeling the drug-target complex accurately, since our simulation involves the modified non-standard Amino Acid, in which the Sotorasib molecule had been glued to the mutated cystein on KRAS. The general process includes the preprocessing of the protein structure, its split into different parts, its format conversion, force field parameters generation, and finally collecting the parts into a complete system ready for the simulation.
Acknowledgements
Author contributions statement
Z. Y., X. Z. and S. Z. conceived and supervised the project. W. L., Z. Y., X. L., D. M., S. Y., Z. Z., C. Z., K. B. and M. D. wrote the code and performed the experiments. W. L., Z. Y., X. L., D. M., S. Y. and Y. C. analyzed the results. W. L., Z. Y., X. L., D. M., S. Y., J. Y., S. Z. wrote the manuscript. All authors reviewed the manuscript.
Data Availability
Code Availability
The code that supports the findings of this study is available at GitHub (https://github.com/AceMapAI/qc-drug-design).
Additional information
The authors declare no competing interest.
Figure 5: (a): Energy evolution during the MD simulation. The energy stabilized after an initial equilibrating phase. (b): Monitoring the QM region energy evolution. (c) The covalent bond is remarkably stable during the whole simulation process. The bond length fluctuates around 1.86 angstrom with a standard deviation less than 0.1 angstrom. The bond length is aligned with previous literature discoveries, and the small deviation is a good indication of the bonding stability.(d) Visualising the bond angle variations during the simulation (CB-SG-C18 and SG-C18-C17).
HF CASCI VQE DFT
$\Delta G^{{\ddagger}}$ with solvent 13.1 11.8 $7\pm 5$ 8.3
$\Delta G$ with solvent -41.1 -49.5 $-51\pm 11$10.8
$\Delta G^{{\ddagger}}$ w/o solvent 2.6 1.4 $-4\pm 7$
$\Delta G$ w/o solvent -59.3 -68.0 $-65\pm 9$